# Multi-messenger signatures of a deformed magnetar in gamma-ray bursts

### v1

@@ 5 GW and EM spin down
This section describes the entire energy loss of the magnetar during EM and GW emissions. Despite the fact that the total luminosity is not an observable quantity by a single instruments, it can illustrate the dynamical behavior of a new born magnetar through multi-messenger emissions.
5.1 Dipole EM and GW Radiations
@@
Here, we consider spin down due to both dipole and GW radiations. The resulting evolutionary equation for eccentricity in incompressible is given by:
@@
and for the compressible model, we have;
@@
Figure 10: Comparing dipole,GW and dipole+GW luminosity for compressible model, Here we assume that $M_{T}=2M\odot$, $B=5\times 10^{15}G$ and $\epsilon=10^{-3}$.
Fig. 10 shows the resulting luminosity from EM dipole and GW radiations for incompressible model of a magnetar with $M_{T}=2M\odot$ and assuming $B=5\times 10^{15}G$ and $\epsilon=10^{-3}$, the total luminosity (EM+GW) is also represented in this figure. According to this plot, it is seen that EM luminosity has a sharper peak which appears with a delay after GW peak time and at the early times GW luminosity is dominated. For a time interval around dipole EM peak, it dominates and then decays with a faster rate than GW luminosity. The total luminosity on the other hand starts to decay earlier and faster with a slope larger than dipole luminosity’s and less than GW one.
It is instructive to compare our results with those presented in Corsi & Mészáros (2009), where it has been shown that the peak amplitude of the GW signal occurs with a delay compare to the GRB trigger, in our scenario we see that for a given set of parameters, the GW flares can also even appear prior to EM ones.
5.2 Dipole and Quadrapole EM and GW Radiations
Taking the quadrapole magnetic moment radiation into our consideration, the evolutionary equations for eccentricity in incompressible and compressible models are given by:
@@ 5.2 Dipole and Quadrapole EM and GW Radiations
Figure 11: GW, EM (dipo+quad) and GW+EM luminosities for the incompressible model assuming $\epsilon=10^{-3}$ and $\kappa=100$ for total mass $M=1.6M\odot$ and $M=2M\odot$ as labeled.
@@
Figure 12: Comparing EM (dipo+quad), GW and total luminosity for a $2M\odot$ magnetar modeled with compressible fluid applying $\Gamma=1.43$ and $\epsilon=10^{-3}$, (top) assuming $B=10^{15}G$ and $\kappa=50$, (bottom) $B=5\times 10^{15}G$, $\kappa=100$.
@@
Fig. 11 illustrates a comparison of the EM (dipo+quad), GW and the total (GW+EM) luminosity for the incompressible model of nascent NSs with $M=1.6M\odot$ and $M=2M\odot$ regarding $\epsilon=10^{-3}$, $\kappa=100$ and $\Gamma=1.43$. We see that the GW luminosity dominates at earlier times and EM radiation overcomes at later times. Similar to the behavior observed in Fig. 10, the total radiation luminosity decays with a faster rate. In Fig. 12, the behavior of EM, GW and total luminosity are illustrated for two different set of parameters. In both the top and bottom panels of Fig. 12, a compressible magnetar with $\Gamma=1.43$ and $\epsilon=10^{-3}$ is assumed, but $B$ and $\kappa$ is different. This figure shows the crucial dependence of the appearance time of EM or GW peaks, in the top panel for $B=10^{15}G$ and $\kappa=50$, the GW peak luminosity occurs perior to the EM peak luminosity, while in the bottom panel the priority of these peaks are changed by assuming $B=5\times 10^{15}G$, $\kappa=100$. Our investigation in this section shows that in the case of appearence of peak within the luminosity there is a delayed arrival time between peak in GW and EM radiations which provide a unique signature can be confirmed by ongoing and upcoming multimessenger observations of high energy astrophysics satellites and ground based detectors.
6 FITting GRB Flares within compressible magnetar model
The GRB afterglow observed in different energy bands from X-ray to the optical and radio bands which follows a similar power-law, but with significant higher luminosity in X-ray band. Consequently, the X-ray luminosity can be served as an indicator of the total (bolometric) afterglow luminosity. It is worth mentioning that due to the anisotropic radiation effects of GRBs introduced by the jet structure and also the instrumental limitations, the bolometric luminosity might be larger than the observed luminosity. Since Swift-XRT generally offers the most comprehensive observations of afterglow evolution, in this paper we rely on the afterglow luminosity data captured by Swift-XRT in order to compare with our model predictions. In some of the previous studies, the bolometric luminosity is estimated by multiplying a factor of 2 - 7 to the energy contained in the Swift-XRT band (0.3-10 keV) Rowlinson et al. (2013); Rueda et al. (2022); Wang et al. (2024).
Here, we have focused on a GRB sample including GRBs 070306, GRB 090812, GRB 100513A and GRB 100901A, which all sources have known redshifts, and their modified light curves presented using k-correction (Bloom et al., 2001). All the GRB sources in our sample have the maximum luminosity of the X-ray flares less than $10^{51}\text{erg s}^{-1}$ and the peak of the luminosity occur in less than 200 s. In order to model the X-ray flares, we consider a new born NS with $2M_{\odot}$ and free parameters are assumed to be the magnetic field $B$, the quadrupole to dipole ratios $\kappa$, and adiabatic index $\Gamma$. Here, we adopt Metropolis-Hasting (M-H) algorithm from Markov Chain Monte Carlo (MCMC) method in Figs. 14 - 16 to find the best fit for deformed newborn magnetar model (presented in sections 2 and 3) taking into account the Swift XRT data. The MCMC python code is also employed in order to generate posterior samples, we plotted a 2d distribution for posteriors in the right pannels of Figs. 14 - 16 showing pairwise correlation between parameters. In order to find the best fit parameters, we used Swift/XRT data for X-flares occurring in less than few hundred seconds (colored purple data points in Figs. 14 - 16) which covers the peak feature of X-ray flares. For each parameters ($B$, $\kappa$ and $\Gamma$), we estimated probability distribution function (PDF) with gaussian kernel density estimation (KDE) on sample’s histogram given in the right pannels of Figs. 14 - 16. The value for width of gaussian function is chosen to be 0.5, showing that the range of valid parameters to fit the data. The dashed magenta lines in PDF plots indicate the median of each parameter sample. We provide $\chi^{2}_{\text{red}}=\chi^{2}/dof$ (chi-squared divided by the degrees of freedom) in Table. 1, and all the fitting parameters summarized in this table with the errors are given as well. The obtained values of fitting parameters shows that $B\sim(6-20)\times 10^{15}G$, $\kappa\sim 49-150$ and $\Gamma\sim 1.5-1.7$ in order to explain the X-ray flares occur in less than few handered seconds after the Swift trigger time. The model presented in this paper is limited to the Maclaurian spheroids, however this investigation can be extended to the Jacobi ellipsoid, in the case that the eccenterity is larger than 0.81 and also for the case in which 3 axial are different.
We find that the peaks in Swift data can be produced by a softer EOS in which $\Gamma$ tends to $1.43$, note that the plateau behaviour in later times is given by stiffer EOSs. Generally magnetic field, $\kappa$ and EOS are not constant thorough dynamical evolution of deformed magnetars, these features should be taken into account in a more realistic configurations, and will be the subject of a future paper. It is worth mentioning to note that while in the present paper, we focus on the modeling of GRB X-ray flares by compressible new born magnetars, our model has the potential to explain the X-ray plateau (in addition to X-ray flares) taking into account different types of the fluids with multipolar structures of the magnetic field Wang et al. (2024); Xie et al. (2022b). We plan to perform a comprehensive statistical analysis for a large sample of X-ray flares in a future paper. Taking into account the time dependency of the magnetic field and also including accretion materials onto a new born magnetar could be origin of the X-ray flares in GRB light curves and can be considered a more generalized model.
7 Conclusions
In this paper, we presented a scenario based on dynamical evolution of a Maclaurin spheroid to model GRB’s central engine taking into account both compressible and incompressible fluids in order to explain different features of EM and GW luminosities. We evaluated the effect of the total mass $M_{T}$, the magnetic field $B$, the ratio of the magnetic quadrupole to dipole moment $\kappa$ and the adiabatic index $\Gamma$ on the light curves of a deformed NS. We showed that the general properties of the EM/GW light curve crucially depends on the nature of EOS which is applied to model the magnetar and we found that for a given set of parameters, our model can generate flaring activities in both GW and EM luminosities. Our main motivation inspired by Finn & Shapiro (1990), where the rotational evolution of a new NS is originating from the deformation of the object which can result in both spin down and spin up motion in various circumstances. Besides considering the EM signal of our GRB model, the GW signal and its detectability in the light of the sensitivity range of GW detectors has been considered. In Sec. 6 for a sample of GRBs by using Swift/XRT data, we applied MCMC method in order to find the best fit parameters of our model. In this study, we tried to provide valuable insights into the underlying mechanisms driving the observed temporal behaviors in GRB afterglows such as the presence of X-ray flares based on a deformed magnetar model, in this regard various multi-messenger predictions of the model have been presented and also the validity of the model has been examined in the light of some of the Swift/XRT observational data.
@@ 7 Conclusions
Figure 13: In the left panel, the compressible model for a magnetar with $M=2M_{\odot}$ is used to fit the Swift/XRT data of GRB 100901A. The colored purple data points (which covers the X-ray flares) are applied to find the best fit model parameters $B_{\text{dip}}$, $\kappa$ and $\Gamma$. In the right panel, the posterior distributions for fitting parameters are plotted taking into account the histogram and estimated PDF for each parameter. The dashed magenta lines in PDF plots illustrate the median of the parameters, and the best fit values are represented at the top the diagrams.
@@
Table 1: Constraints on properties of the deformed newborn magnetar via fitting the X-ray light curve of GRBs. The table presents the fitting results, including the magnetic field strength (B), the adiabatic index $\Gamma$, and the ratio of the magnetic quadrupole to dipole moment ($\kappa$).
@@ Acknowledgments
S.S and P.H would appreciate for valuable discussions during the conferences of High Energy Astrophysics and Cosmology in the era of all-sky in October 2024, Yerevan, Armenia. S.S and P.H are grateful to Massimo Della Valle for insightful comments and discussions.
@@
Stability analysis and Black Hole formation
In this appendix, we consider the criteria of the black hole (BH) formation for both incompressible and compressible models in order to determine allowed region of the parameter space over the mass and adiabatic index. For the incompressible model, we should evaluate the Schwarzschild radius at $a_{0}=a(e=0)$ to investigate whether the BH is formed or not. But in the case of compressible model, semi major axis $a$ is a more complicated function of the eccentricity and it is not obvious that if $a_{0}$ gives the minimum radius. Taking into account the logarithm of Eq. (11), we find that :
@@
where $a_{\text{crit}}\sim R_{S}=2M$. Therefore, one can find a critical value for $\Gamma$ as (see Eq. 63 Finn & Shapiro (1990))
@@
For a NS with mass M if $\Gamma<\Gamma_{\text{crit}}$ at a given value of eccentricity, the star will collapse to a BH eventually. The stability analysis of the NS are displayed in Figs. 17 and 18. In left panel of Fig. 17, we present $\Gamma_{\text{crit}}$ for various masses in range $(1.6-2.6)M_{\odot}$ as a function of eccentricity. For all chosen masses, all $\Gamma_{\text{crit}}$ curves lies below $\Gamma=1.43$ line at the given range of eccentricity meaning that a NS with a mass within this range remains stable during its evolution. On the right panel of Fig. 17, the mass range $(3.6-3.9)M_{\odot}$ is adopted, within this range all $\Gamma_{\text{crit}}$ curves are above $\Gamma=1.43$ line at the beginning of evolution (for larger e) and they cross the horizontal line at some eccentricity values. This means that for a range of M and e, we can not have stable configuration for $\Gamma=1.43$ and therefore this part of parameter space is excluded in our consideration.
In Fig. 18, we present a density plot for $\Gamma_{\text{crit}}$ as a function of mass and eccentricity. The presence of a singularity in the denominator of Eq. (A3) for some values of mass and eccentricity leads to infinite value for $\Gamma_{\text{crit}}$, showing in the plot with the sharper color. Regarding the discussion mentioned above, this plot provide essential information about the maximum allowed value of $\Gamma$ and therefore the stable region of our parameter space.
Figure 17: The plot shows the variation of the adiabatic index for various masses as labeled, The colored lines show $\Gamma_{\text{crit}}$ for different masses and the horizontal dashed lines display $\Gamma=1.43$.
Figure 18: The density plot $\Gamma_{\text{crit}}$ is displayed as a function of both mass and eccentricity. This figure contains essential information about the stability of NS which is plotted in log scale.

### v2

@@ 5 GW and EM spin-down
This section describes the entire energy loss of the magnetar during EM and GW emissions. Despite the fact that the total luminosity is not an observable quantity by a single instrument, it can illustrate the dynamical behavior of a new born magnetar through multimessenger emissions.
5.1 Dipole EM and GW Radiation
@@
Here, we consider spin down due to both dipole and GW radiations. The resulting evolutionary equation for eccentricity in incompressible is given by
@@
and for the compressible model, we have
@@
Figure 10: Comparing dipole, GW, and dipole+GW luminosity for the compressible model. Here, we assume that $M_{T}=2M\odot$, $B=5\times 10^{15}G$ and $\epsilon=10^{-3}$.
Fig. 10 shows the resulting luminosity from EM dipole and GW radiations for incompressible model of a magnetar with $M_{T}=2M\odot$ and assuming $B=5\times 10^{15}G$ and $\epsilon=10^{-3}$, the total luminosity (EM+GW) is also represented in this figure. According to this plot, it is seen that EM luminosity has a sharper peak, which appears with a delay after the GW peak time, and at the early times the GW luminosity, is dominated. For a time interval around dipole EM peak, it dominates and then decays with a faster rate than GW luminosity. The total luminosity, on the other hand, starts to decay earlier and faster with a slope larger than the dipole luminosity and less than GW one.
It is instructive to compare our results with those presented in Corsi & Mészáros (2009), where it has been shown that the peak amplitude of the GW signal occurs with a delay compared to the GRB trigger. In our scenario, we see that for a given set of parameters, the GW flares can also even appear prior to EM ones.
5.2 Dipole and Quadrupole EM and GW Radiation
Taking the quadrupole magnetic moment radiation into our consideration, the evolutionary equations for eccentricity in the incompressible and compressible models are given by
@@ 5.2 Dipole and Quadrupole EM and GW Radiation
Figure 11: GW, EM , and GW+EM luminosities for the incompressible model assuming $\epsilon=10^{-3}$ and $\kappa=100$ for total mass $M=1.6M\odot$ and $M=2M\odot$ as labeled.
@@
Figure 12: Comparing EM (dipo+quad), GW, and total luminosity for a $2M\odot$ magnetar modeled with compressible fluid, applying $\Gamma=1.43$ and $\epsilon=10^{-3}$, (top) assuming $B=10^{15}G$ and $\kappa=50$, (bottom) $B=5\times 10^{15}G$, $\kappa=100$.
@@
Fig. 11 illustrates a comparison of the EM, GW, and the total (GW+EM) luminosity for the incompressible model of nascent NSs with $M=1.6M_{\odot}$ and $M=2M_{\odot}$ regarding $\epsilon=10^{-3}$, $\kappa=100$ and $\Gamma=1.43$. We see that the GW luminosity dominates at earlier times and EM radiation overcomes at later times. Similar to the behavior observed in Fig. 10, the total radiation luminosity decays at a faster rate. In Fig. 12, the behavior of EM, GW, and total luminosity are illustrated for two different sets of parameters. In both the top and bottom panels of Fig. 12, a compressible magnetar with $\Gamma=1.43$ and $\epsilon=10^{-3}$ is assumed, but $B$ and $\kappa$ are different. This figure shows the crucial dependence of the appearance time of EM or GW peaks. In the top panel for $B=10^{15}G$ and $\kappa=50$, the GW peak luminosity occurs prior to the EM peak luminosity, while in the bottom panel, the priority of these peaks are changed by assuming $B=5\times 10^{15}G$, $\kappa=100$. Our investigation in this section shows that in the case of the appearance of a peak within the luminosity, there is a delayed arrival time between peak in the GW and EM radiation, which provides a unique signature that can be confirmed by ongoing and upcoming multimessenger observations of high-energy astrophysics satellites and ground-based detectors.
@@
Figure 13: In the left panel, the compressible model for a magnetar with $M=2M_{\odot}$ is used to fit the Swift-XRT data of GRB 100901A. The colored purple data points (which cover the X-ray flares) are applied to find the best-fit model parameters $B_{\text{dip}}$, $\kappa$ and $\Gamma$. In the right panel, the posterior distributions for fitting parameters are plotted taking into account the histogram and estimated the PDF for each parameter. The dashed magenta lines in PDF plots illustrate the median of the parameters, and the best-fit values are represented in the top of the diagrams.
@@
6 Fitting GRB Flares within the compressible Magnetar Model
The GRB afterglow is observed in different energy bands from X-ray to the optical and radio bands, which follows a similar power law, but with significantly higher luminosity in the X-ray band. Consequently, the X-ray luminosity can serve as an indicator of the total (bolometric) afterglow luminosity. It is worth mentioning that due to the anisotropic radiation effects of GRBs introduced by the jet structure and also the instrumental limitations, the bolometric luminosity might be larger than the observed luminosity. Since the Swift-X-ray Telescope (XRT) generally offers the most comprehensive observations of afterglow evolution, in this paper, we rely on the afterglow luminosity data captured by Swift-XRT in order to compare with our model predictions. In some of the previous studies, the bolometric luminosity is estimated by multiplying a factor of 2 - 7 to the energy contained in the Swift-XRT band (0.3-10 keV) Rowlinson et al. (2013); Rueda et al. (2022); Wang et al. (2024).
Here, we have focused on a GRB sample, including GRB 070306, GRB 090812, GRB 100513A and GRB 100901A, of which have known redshifts, and their modified light curves presented using a k-correction (Bloom et al., 2001). All the GRB sources in our sample have the maximum luminosity of the X-ray flares less than $10^{51}\text{erg s}^{-1}$ and the peak of the luminosity occurs in less than 200 s. In order to model the X-ray flares, we consider a new born NS with $2M_{\odot}$ and free parameters are assumed to be the magnetic field $B$, the quadrupole-to-dipole ratio $\kappa$, and the adiabatic index $\Gamma$. Here, we adopt the Metropolis-Hasting (M-H) algorithm from the Markov Chain Monte Carlo (MCMC) method in Figs. 13 - 16 to find the best fit for the deformed newborn magnetar model (presented in sections 2 and 3) taking into account the Swift-XRT data. The MCMC Python code is also employed in order to generate posterior samples. We plotted a 2D distribution for posteriors in the right panels of Figs. 13 - 16 showing a pairwise correlation between parameters. In order to find the best-fit parameters, we used Swift-XRT data for X-flares occurring in less than a few hundred seconds (colored purple data points in Figs. 13 - 16) which covers the peak feature of the X-ray flares. For each parameter ($B$, $\kappa$ and $\Gamma$), we estimated the probability distribution function (PDF) with a Gaussian kernel density estimation (KDE) on the sample’s histogram given in the right panels of Figs. 13 - 16. The value for the width of Gaussian function is chosen to be 0.5, showing that the range of valid parameters to fit the data. The dashed magenta lines in the PDF plots indicate the median of each parameter sample. We provide $\chi^{2}_{\text{red}}=\chi^{2}/dof$ (chi-squared divided by the degrees of freedom) in Table. 2, and all the fitting parameters summarized in this table, with the errors are given as well. The model presented in this paper is limited to the stable Maclaurin spheroids; however, this investigation can be extended to the Jacobi ellipsoid, in the case that the eccentricity is larger than 0.81 and also for the case in which the three axes are different. Another direction to follow is considering spheroids in the presence of perturbations, taking into account the parameter space which is open to CFS instabilities, and to examine multimessenger signals in this regime.
We find that the peaks in Swift data can be produced by a softer EoS in which $\Gamma$ tends to $1.43$. Note that the plateau behavior in later times is given by stiffer EoSs. Generally, the magnetic field, $\kappa$ and EoS are not constant throughout the dynamical evolution of the deformed magnetars. These features should be taken into account in a more realistic configuration and will be the subject of a future paper. It is worth mentioning to note that while in the present paper, we focus on the modeling of GRB X-ray flares by compressible new born magnetars, our model has also the potential to explain the X-ray plateau (in addition to X-ray flares) taking into account different types of the fluids with multipolar structures of the magnetic field Wang et al. (2024); Xie et al. (2022b). We plan to perform a comprehensive statistical analysis for a large sample of X-ray flares in a future paper. Taking into account the time dependency of the magnetic field and also including accretion materials onto a new born magnetar could be the origin of the X-ray flares in GRB light curves and can be considered in a more generalized model.
7 Conclusions
In this paper, we presented a scenario based on the dynamical evolution of a Maclaurin spheroid to model the GRB’s central engine, taking into account both compressible and incompressible fluids in order to explain different features of EM and GW luminosities. We evaluated the effect of the total mass $M_{T}$, the magnetic field $B$, the ratio of the magnetic quadrupole-to-dipole moment $\kappa$ and the adiabatic index $\Gamma$ on the light curves of a deformed NS. As it was presented in Sec. 6, our model was applied to fit X-ray flares in a GRB sample observed by Swift-XRT with the peak luminosity $\sim 10^{49}$ $10^{50}$ $\rm\text{erg}s^{-1}$ and the peak time less than $200$ s. The obtained values of the fitting parameters are $B\sim(6-20)\times 10^{15}G$, $\kappa\sim 49-150$ and $\Gamma\sim 1.5-1.7$. Moreover, considering a reasonable range of parameters, our model has the potential to explain X-ray flares with the peak luminosity on the order of $10^{46}$ $10^{51}$ $\rm\text{erg}/s$ which occur within the wide range of $10$ to $10^{4}$ seconds after the Swift trigger time.
Our main motivation is inspired by Finn & Shapiro (1990), where the rotational evolution of a new NS originates from its deformation, which can result in both spin-down/spin-up motion in various circumstances. We extended the analysis presented in Finn & Shapiro (1990) to be used for the GRB central engine. In addition to performing the stability analysis over the parameter space, we introduced an additional quadrupole magnetic radiation term, which showed that the temporal spin-up of a newborn magnetar, for some specific EoSs, can lead to flaring features in the light curves. Previous studies primarily focused on explaining the X-ray plateau in the EM light curves of GRBs based on the rotational energy loss of magnetars Wang et al. (2024); Xie et al. (2022b); Rueda et al. (2022); Lasky & Glampedakis (2016). However, in our work, we account for rotational distortion induced by the rapid rotation of NSs, which causes angular momentum to change due to the star’s deformation. In our approach, the spin variation of a newborn magnetar is derived from the equilibrium equations of the Maclaurin spheroid, where the internal structure of the system through various EoSs affects the spin behavior of NSs.
In addition to analyzing the EM signal in our GRB model, we have examined the GW signal and its detectability based on the sensitivity range of GW detectors, as discussed in Sec. 4. Using a reasonable set of parameters, the GW luminosity predicted by the model is $\sim 10^{48}$ $10^{52}$ $\rm\text{erg}/s$. Meanwhile, the peak time crucially depends on the model parameters and applied EoS, and the delay time between the peaks in GW and EM radiation (see Fig. (12)) can serve as a unique signature for current and future multimessenger observations. By evaluating the characteristic GW strain for both spin-down regimes driven by EM and GW radiation, we have considered the detectability of the GW signals generated by a newly born deformed magnetar. A peculiar signature in the GW strain, tracing the spin evolution of the magnetar and its EoS, is also predicted in Fig. 9. The SNR for different GW detectors - ALIGO O4, ALIGO O5, KAGRA O5, ET, and CE - is estimated for a central frequency of 1 kHz in Table. 1. This analysis shows that ET and CE have the highest detectability potential for magnetar GW signals.
In summary, through this study, we aimed to provide valuable insights into the underlying mechanisms driving the observed temporal behaviors in early-time GRB afterglows, such as the presence of X-ray flares, by considering a deformed magnetar as the GRB central engine. We demonstrated that the general properties of EM and GW luminosities crucially depend on the nature of the magnetar’s EoS, which can generate flaring activities in both signals. Additionally, we predicted several distinct multimessenger features, highlighting the great potential of upcoming EM and GW missions to shed light on the nature of the GRB central engine.
@@ 7 Conclusions
Table 2: Constraints on the properties of the deformed newborn magnetar via fitting the X-ray light curve of GRBs. The table presents the fitting results, including the magnetic field strength (B), the adiabatic index $\Gamma$, and the ratio of the magnetic quadrupole-to-dipole moment ($\kappa$).
@@ Acknowledgments
We thank the anonymous referee for the valuable comments and suggestions. S.S and P.H would appreciate for valuable discussions during the conferences of High Energy Astrophysics and Cosmology in the era of all-sky in October 2024, Yerevan, Armenia. S.S and P.H are grateful to Massimo Della Valle for insightful comments and discussions.
@@
Stability analysis and BH formation
In this appendix, we consider the criteria of the BH formation for both the incompressible and compressible models in order to determine allowed region of the parameter space over the mass and adiabatic index. For the incompressible model, we should evaluate the Schwarzschild radius at $a_{0}=a(e=0)$ to investigate whether the BH is formed or not. But in the case of the compressible model, semi major axis $a$ is a more complicated function of the eccentricity, and it is not obvious that if $a_{0}$ gives the minimum radius. Taking into account the logarithm of Eq. (12), we find that
@@
where $a_{\text{crit}}\sim R_{S}=2M$. Therefore, one can find a critical value for $\Gamma$ as (see Eq. 63) inFinn & Shapiro (1990)
@@
For an NS with mass M, if $\Gamma<\Gamma_{\text{crit}}$ at a given value of eccentricity, the star will collapse to a BH eventually. The stability analysis of the NS are displayed in Figs. 17 and 18. In the left panel of Fig. 17, we present $\Gamma_{\text{crit}}$ for various masses in the range $(1.6-2.6)M_{\odot}$ as a function of eccentricity. For all chosen masses, all $\Gamma_{\text{crit}}$ curves lie below the $\Gamma=1.43$ line at the given range of eccentricity, meaning that an NS with a mass within this range remains stable during its evolution. On the right panel of Fig. 17, the mass range $(3.6-3.9)M_{\odot}$ is adopted. Within this range, all $\Gamma_{\text{crit}}$ curves are above $\Gamma=1.43$ line at the beginning of the evolution (for larger e), and they cross the horizontal line at various eccentricity values. This means that for a range of M and e, we are unable to have a stable configuration for $\Gamma=1.43$ and therefore, this part of the parameter space is excluded from our consideration.
In Fig. 18, we present a density plot for $\Gamma_{\text{crit}}$ as a function of mass and eccentricity. The presence of a singularity in the denominator of Eq. (A3) for some values of mass and eccentricity leads to an infinite value for $\Gamma_{\text{crit}}$, showing in the plot with the sharper color. Regarding the discussion mentioned above, this plot provides essential information about the maximum allowed value of $\Gamma$ and therefore, the stable region of our parameter space.
Figure 17: The plot shows the variation of the adiabatic index for various masses as labele. The colored lines show $\Gamma_{\text{crit}}$ for different masses, and the horizontal dashed lines display $\Gamma=1.43$.
Figure 18: The density plot $\Gamma_{\text{crit}}$ is displayed as a function of both mass and eccentricity. This figure contains essential information about the stability of NS, which is plotted in log scale.